OpenStudy (anonymous):

HELP!! simplify? 7/8(x+16)-1/4(32-2x) i dont understand this one!

OpenStudy (lilg132):

7/8 = 0.875 so use that instead of the fraction it should make it easier

OpenStudy (lilg132):

-1/4 = -0.25

OpenStudy (anonymous):

ok hold on let me try it

OpenStudy (anonymous):

ok so far i have

OpenStudy (anonymous):

0.875 * 16 = 14?

OpenStudy (anonymous):

then -0.25*32=-8

OpenStudy (anonymous):

-0.25*2=-0.5x?

OpenStudy (anonymous):

1/8 (11 x+48)

OpenStudy (anonymous):

then 0.875x+-0.5x =0.375?

OpenStudy (anonymous):

0.375 + 6?

OpenStudy (anonymous):

@lilg132 help?

OpenStudy (lilg132):

sorry was answering another question

OpenStudy (anonymous):

its ok

OpenStudy (lilg132):

ok let me do it in steps again i havent really checked if your right but you can check with my answers 7/8 = 0.875 -1/4 = -0.25 7/8(x+16)-1/4(32-2x) replace the fractions with the numbers so you would have 0.875(x + 16) -0.25(32 - 2x) 0.875 * x = 0.875x 0.875 * 16 = 14 -0.25 * 32 = -8 -0.25 * -2x = 0.5x (REMEMBER - * - ALWAYS GIVES +) ok so now we have 0.875x + 0.5x + 14 - 8 which gives you 1.375x + 6

OpenStudy (anonymous):

i dont need to change that to a fraction tho?

OpenStudy (lilg132):

not if you find it easier with fractions

OpenStudy (lilg132):

its just my way of doing it i find it easier working with numbers rather than fractions in the end you will always get the same answer and you can change the numbers back to fractions

OpenStudy (anonymous):

oh ok, i rather do it like this i dont really understand fractions :(

OpenStudy (lilg132):

1 3/8x + 6

OpenStudy (anonymous):

13 over 8?

OpenStudy (lilg132):

no 1 and 3/8x

OpenStudy (anonymous):

oh okay

OpenStudy (lilg132):

but dont worry about that so much as long as you have the correct number the answer is fine

OpenStudy (anonymous):

okay then!

OpenStudy (lilg132):

and you done most of the steps correct except just always remember - * - = + - * + = - + * - = - + * + = +

OpenStudy (anonymous):

yes lol the neg and pos always confuse me /:

OpenStudy (lilg132):

well im off now see you next time bye